Water waves in a lake travel 4.73 m in 1.01 s. The period of oscillation is 1 s. What is the speed of the water waves? Answer in units of m/s.
1
Expert's answer
2011-02-01T04:37:43-0500
The speed of the wave is equal to
v = 4.73/1.01 = 4.68 m/s
The frequency is f = 1/T = 1/1[s] = 1 [s-1]
The wavelength whould be
λ = v/f = 4.68 m.